A company uses activity-based costing to determine the costs of its three products: A, B, and C. The budgeted cost and activity for each of the company's three activity cost pools are shown in the following table:

Budgeted Activity
Activity Cost Pool Budgeted Cost Product A Product B Product C
Activity 1 $ 85,000 7,500 10,500 21,500
Activity 2 $ 60,000 8,500 16,500 9,500
Activity 3 $ 112,000 4,000 2,500 3,125


How much overhead will be assigned to Product B using activity-based costing?

Multiple Choice

  • $257,000.00

  • $99,151.20

  • $85,000.00

  • $75,241.11

  • $80,381.50
Answer #1

E. $80,381.50

Activity 1: $85,000/(7,500 + 10,500 + 21,500) = $2.1519

Activity 2: $60,000/(8,500 + 16,500 + 9,500) = $1.7391

Activity 3: $112,000/(4,000 + 2,500 + 3,125) = $11.6364

Overhead assigned to Product B using activity-based costing = (10,500)( $2.1519) + (16,500)($1.7391) + (2,500)($11.6364) = $80,381.50